OpenStudy (anonymous):

solving equations using the zero property and the problem is z^2 - 10z = -21 (what do I do)

OpenStudy (mertsj):

Set equal to 0 and factor.

OpenStudy (mertsj):

z^2-10z+21=0 (z-7)(z-3)=0 So z-7=0 or z-3=0 z=7 or z=3

OpenStudy (anonymous):

thank you I did it right then!!!

OpenStudy (anonymous):

how do you do y^2 - 28 = 0 using the square root property of equality